please help me solve this problem.
Solve the equation: t = 2 + t – 3
4 3

Answer by Alan3354(69443)

Solve the equation: t = 2 + t – 3
t/4 = (t-3)/3
Multiply by 12
3t = 4*(t-3)
3t = 4t - 12
-t = -12
t = 12